If each edge of a cube is doubled, 
how many times will its surface area increase?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the edge of cube be“l”.
Formula for Surface area of the cube, A = 6 l^2
When edge of cube is doubled, then
Surface areaof the cube, say A’ = 6(2l)^2= 6×4l^2= 4(6 l^2)A’ = 4A
Hence, surface area will increase by four times.

Thanks